OpenStudy (anonymous):

Write a radical in which the radicand consists of a coefficient and a variable. Use a coefficient that can be simplified best using the greatest perfect square method. Then, simplify the radical completely. You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a).

jimthompson5910 (jim_thompson5910):

did you manage to get anywhere with this?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

I made this \[\sqrt{18x}\] now after this I don't know what to do

jimthompson5910 (jim_thompson5910):

ok good start

OpenStudy (anonymous):

what do I do next

jimthompson5910 (jim_thompson5910):

can you factor 18 in such a way where one of the factors is a perfect square?

OpenStudy (anonymous):

yeahh 3

OpenStudy (anonymous):

that make 9*2=18

jimthompson5910 (jim_thompson5910):

good, 18 = 9*2 this means that \[\large \sqrt{18a} = \sqrt{9*2a}\] \[\large \sqrt{18a} = \sqrt{9}*\sqrt{2a}\] \[\large \sqrt{18a} = 3\sqrt{2a}\]

OpenStudy (anonymous):

thnk u

jimthompson5910 (jim_thompson5910):

you're welcome